### v1

# Boosting Open Set Recognition Performance through Modulated Representation Learning

Abstract
The open set recognition (OSR) problem aims to identify test samples from novel semantic classes that are not part of the training classes, a task that is crucial in many practical scenarios. However, existing OSR methods use a constant scaling factor (the temperature) to the logits before applying a loss function, which hinders the model from exploring both ends of the spectrum in representation learning – from instance-level to semantic-level features. In this paper, we address this problem by enabling temperature-modulated representation learning using our novel negative cosine scheduling scheme. Our scheduling lets the model form a coarse decision boundary at the beginning of training by focusing on fewer neighbors, and gradually prioritizes more neighbors to smooth out rough edges. This gradual task switching leads to a richer and more generalizable representation space. While other OSR methods benefit by including regularization or auxiliary negative samples, such as with mix-up, thereby adding a significant computational overhead, our scheme can be folded into any existing OSR method with no overhead. We implement the proposed scheme on top of a number of baselines, using both cross-entropy and contrastive loss functions as well as a few other OSR methods, and find that our scheme boosts both the OSR performance and the closed set performance in most cases, especially on the tougher semantic shift benchmarks [^1] The project codes are available at https://anonymous.4open.science/r/NegCosSch-4516/..
(a) $\tau=0.5$
(b) $\tau=1.0$
(c) $\tau=2.0$
(d) CosSch
(e) NegCosSch – starting of the final period
(f) NegCosSch – as training progresses
(g) NegCosSch – when training finishes
Figure 1: UMAP projection of representation spaces for different temperature schedules on 10 classes of the Caltech-UCSD-Birds dataset. (a) - (c) show representations for constant temperatures ($\tau$). For lower $\tau$, the representations of open set samples are distributed and so are the representations of closed set classes, leading to a sharp decision boundary. For higher $\tau$, the representations of the known classes are more compact, making the decision boundary smoother. However, open set samples overlap with the clusters of known classes. Mid value of $\tau$ achieves a trade-off but does not gain the benefits of both ends. (e) - (g) show representations of our temperature scheduling scheme NegCosSch as the training progresses. A lower $\tau$ at the start leads to a coarse decision and the model gradually makes the classes more compact and the open set representations are pushed away. Finally, (d) show the representation space for a previous temperature scheduling scheme CosSch, which is better than constant temperatures but not as compact as our NegCosSch. Experimental details appear in the Appendix.
1 Introduction
Deep learning models have shown impressive performance by learning useful representations particularly for tasks involving the classification of examples into categories present in the training dataset, also known as the closed set. However during inference, in many practical scenarios, test samples may appear from unknown classes (termed as the open set), which were not a part of the training set. Hence, a more realistic task known as the open set recognition (OSR) (Scheirer et al. (2012); Chen et al. (2020a)) aims to simultaneously flag the test samples from unknown classes while accurately classifying examples from the known classes, requiring strong generalization beyond the support of training data.
Early research attempts model the unknown classes as a long-tailed distribution using extreme value theory (Vignotto and Engelke (2018)) to modify the softmax function (Bendale and Boult (2016)), while other methods generate unknown samples either with mix-up or using generative models to represent novel classes (Zhou et al. (2021); Ge et al. (2017); Neal et al. (2018); Chen et al. (2021)). The synthetic examples may not generalize well to a variety of unknown classes. Later methods add regularization to explicitly bound the risks for novel classes (Zhou et al. (2021); Chen et al. (2021, 2020a)). In essence, these methods create more empty regions in the representation space hoping that unknown representations lie in those regions. The tougher unknown samples lie very close to the known classes in representation space, which reduce the effectiveness of such methods.
Vaze et al. (2022) establish new OSR baselines by training models with better design choices and argue that well-trained closed set classifiers achieve better OSR performance, where the unknown samples exhibit lower max logit scores. This essentially has motivated the next generation of OSR methods to learn even better representations for improving OSR performance. The recent methods essentially aim for better representation learning with a better loss function, such as the contrastive loss (Khosla et al. (2020); Chen et al. (2020b)) with sample mix-up (Verma et al. (2018); Zhang et al. (2017)) and by adding different regularization schemes (Zhou et al. (2024); Bahavan et al. (2025); Li et al. (2025); Wang et al. (2025)).
Moreover, the regular OSR benchmarks commonly used are small in scale. In this regard, semantic shift benchmarks (SSBs) are proposed by Vaze et al. (2022) on three fine-grained datasets, having more classes with varying levels of OSR difficulty. Therefore, the OSR methods that demonstrate improvement on smaller datasets but involve either data generation (Chen et al. (2021); Ge et al. (2017)), mix-up (Zhou et al. (2021); Xu et al. (2023)), or training an additional model with separate objective, such as VAEs that include reconstruction based objective (Oza and Patel (2019); Yoshihashi et al. (2019); Zhou et al. (2024)), are computationally demanding and often require higher memory. These methods are unsuitable for SSBs as the base model for SSBs requires a significant compute for good closed set performance. Most of the latest research does not use these benchmarks. Therefore, better representation learning schemes are required, which add little to no computational overhead.
To achieve this, we need to explore the inner mechanisms of the losses that are the basis of most OSR methods, such as the cross-entropy (CE) loss and the contrastive loss. These loss functions compute probabilities by applying a temperature scaling to the logits- the model’s raw outputs- where the temperature coefficient adjusts the sharpness of the resulting probability distributions. Temperature is the key parameter to control the learned features for both losses. Works by Zhang et al. (2022); Kukleva et al. (2023); Zhang et al. (2021) demonstrate that learning instance-specific representations are encouraged by a lower temperature value, while semantic representations are learned through a higher value. However, a fixed temperature throughout the training prevents the model from exploring both ends of this learning spectrum. In this regard, Kukleva et al. (2023) demonstrate the benefits of learning both instance-level and semantic features in self-supervised training for long-tailed datasets in closed set scenarios using a cosine temperature scheduling.
In OSR, learning a representation space that provides both instance-specific and semantic features is also crucial to achieve better open set and closed set performance. In this research, we analyze the representation space for different temperature values on both losses. Based on our analysis, we formulate a novel temperature scheduling scheme for temperature modulated representation learning. We find temperature modulation with the proposed scheduling is beneficial to create more compact clusters for representing closed set classes, while keeping more open set examples distant from these clusters, resulting in overall better representations.
The main contributions of this paper are summarized as follows:
  • We propose a novel temperature scheduling (TS) scheme called the negative cosine scheduling (NegCosSch) for temperature modulated representation learning. The scheme can be seamlessly integrated into any existing OSR algorithm, with both the CE and the losses based on contrastive learning, without any computational overhead. We show performance improvements on both the regular OSR benchmarks and the SSBs.
  • We further integrate our method with a few existing methods for OSR and show a significant boost in their performance.
  • Our scheduling demonstrates strong performance improvements for the tougher SSBs over the baselines on both the closed set and the open set problems. We show that our scheme brings out stronger improvements with an increased number of training classes when the task becomes more difficult for the baseline model.
The rest of the paper is organized as follows. In Section 2, we discuss the relevant background with different losses. In Section 3, we describe the proposed TS scheme. In Section 4, we discuss our results followed by the related works in Section 5 and the concluding remarks in Section 6.
2 Background
During training, the model is decomposed into two components: The first component is an encoder function $f(\cdot)$ which maps the input $x$ to a representation $z=f(x)$. The second component $h(\cdot)$ maps the representations to task specific outputs, which is either a linear classification layer if we train with the CE loss or a projection layer if we use the contrastive training. The final outputs, also called the logits, $l=h(z)=h(f(z))$ are then given as model predictions to the loss functions. We assume for a specific problem, a model is trained for a predefined number of epochs $E$. We further discuss the CE and the supervised contrastive (SupCon) losses and the effects of the temperature parameter, which provide a basis for our proposed scheme.
2.1 Cross-entropy Loss
For a batch of training data =$\mathcal{B}=\{(x_{k},y_{k})\}_{k=1}^{B}$, the CE loss is calculated as
$L_{\text{CE}}=-\frac{1}{|\mathcal{B}|}\sum_{(x_{k},y_{k})\in\mathcal{B}}\text{one\_hot}(y_{k})\cdot\text{log}(p_{k})$ (1)
[…]
2.2 Supervised Contrastive Loss
For a given batch $\mathcal{B}$, the SupCon training utilizes a multi-viewed batch by taking two augmented samples of the same original sample. The multi-viewed batch =$\mathcal{B}^{\prime}=\{(\tilde{x_{i}},\tilde{y}_{i})\}_{i=1}^{2B}$, where $\tilde{x}_{2k}$ and $\tilde{x}_{2k-1}$ are two random augmentations of $x_{k}~{}~{}(k=1,...,B)$, and $\tilde{y}_{2k-1}=\tilde{y}_{2k}=y_{k}$. Below we refer to $i$ as the anchor index from $I=\{1,...,2B\}$. The SupCon loss (Khosla et al. (2020)) is defined by:
$L_{\text{SupCon}}=-\frac{1}{|I|}\sum_{i\in I}\frac{1}{|P(i)|}\Bigg{[}\sum_{p\in P(i)}\text{log}\frac{\text{exp}(\text{sim}(l_{i},l_{p})/\tau)}{\sum_{a\in I\setminus\{i\}}\text{exp}(\text{sim}(l_{i},l_{a})/\tau)}\Bigg{]}$ (2)
Here, $\text{sim}(l_{i},l_{j})$ is the cosine similarity between $l_{i}$ and $l_{j}$. $P(i)=\{p\neq i:\tilde{y}_{i}=\tilde{y}_{p}\}$ is the set of indices of all samples having the same label as $\tilde{y}_{i}$ in $\mathcal{B}^{\prime}$ distinct from $i$. Contrastive loss, by construction, gains its strength by pushing away the representations of the negative samples (samples of other classes) and by producing compact clusters of representations for the (positive) samples of the same class.
2.3 Effect of Temperature on the Losses
The SupCon loss applies hard negative mining from penalizing the harder negative samples more through the exponential function (Khosla et al. (2020)). The measure of hardness of a sample with respect to an anchor sample is determined by the similarity function. Therefore as a scaling factor, the temperature plays a critical role in controlling the trade-off between uniformity and semantic structure in the representation space as shown in Wang and Liu (2021); Kukleva et al. (2023) for the self-supervised loss (Chen et al. (2020b)). This effect mostly translates to the supervised case except for the fact that the definition of positive and negative samples are now different.
For smaller values of temperature $\tau$, as the differences in similarity get amplified, the nearest negative samples receive the highest gradient (Wang and Liu (2021)). The model aggressively pushes the negative samples away, leading to features that are appropriate for instance-level discrimination and distributing the embeddings over the representation space. However, the positive samples do not cluster tightly because, like the negative samples, fewer positive neighbors get priority in the loss function (Figure 1(a)). The resulting decision boundary is sharper. The open set samples do not get closer to the known classes due to the heavy penalty of having slight dissimilarity.
With larger $\tau$, the differences in similarity diminish and the repulsive force gets distributed to more negative neighbors. The model can decrease the loss by learning the group-wise features rather than the instance discriminating features to push away easy negatives, inducing semantic structures. Due to compact clusters of within-class representations, the resulting decision boundary is smoother. However, the between-class distances also reduce as the model is less aggressive in removing the negatives. In this case a lot of open set examples get close to the known classes (Figure 1(c)).
Similarly in CE loss, lower values of temperature ($\tau<1$) leads to a sharper output probability distribution over the training classes (Guo et al. (2017)), while the higher values of $\tau>1$ makes the output probability distribution smoother.
The value of $\tau$ is usually kept constant throughout the entire training for both losses, which is set either to a predefined value or chosen with hyperparameter tuning.
3 Proposed Method
In this section, we formally introduce our problem, describe our proposed TS scheme and how the scheduling leads to learning representations useful for OSR.
3.1 Problem Definition
We are given a labeled training dataset $\mathcal{D}_{tr}=\{(x_{i},y_{i})\}\subset\mathcal{X}\times\mathcal{Y}$, where $x_{i}$ is the training sample with label $y_{i}$. $\mathcal{X}$ is the input space and the labels of $\mathcal{D}_{tr}$ from a closed set label space $\mathcal{Y}$, i.e. $y_{i}\in\mathcal{Y},\forall i$. The total number of classes in the closed set is $C=|\mathcal{Y}|$. The test dataset $\mathcal{D}_{test}\subset\mathcal{X}\times\mathcal{Y}\cup\mathcal{O}$ consists of samples whose label space $\mathcal{Y}\cup\mathcal{O}$ is different than $\mathcal{Y}$, and $\mathcal{O}\cap\mathcal{Y}=\emptyset$. $\mathcal{O}$ is the set of unknown classes termed as the open set. The objective of OSR is to classify a test sample among the closed set classes or to flag it as belonging to an unknown class. We assume that information about the nature of unknown classes or any auxiliary samples are unavailable during training.
Figure 2: Different temperature scheduling schemes.
3.2 Proposed Temperature Scheduling
For a temperature dependent loss function, $\tau$ is usually kept constant throughout the training process. We denote constant TS by Const, defined as
$\text{Const}(e;\tau)=\tau,\forall e$ (3)
Where $e$ is the epoch number.
For OSR, if the test representation from a novel class is too class-specific, the model easily finds its similarity to one of the known classes. On the other hand, if the feature is too instance-specific, the model is under-confident in assigning any sample to a known class. Therefore, a model needs to capture both features – good class-specific representations having room for instance level discriminating power within the class. A specific temperature throughout the training restricts the opportunity to explore this trade-off that may lead to learning better representations. We propose to gradually switch between the two objectives using a generalized cosine scheduling, which we describe next.
3.2.1 Generalized Cosine Scheduling
Instead of a constant value, we propose to schedule the temperature from a range $[\tau^{-},\tau^{+}]$ as the training progresses using a generalized cosine scheduling, which is defined as
$\text{GCosSch}(e;\tau^{+},\tau^{-},P,k)=\left\{\begin{array}[]{ll}\tau^{-}+\frac{1}{2}(\tau^{+}-\tau^{-})(1+\text{cos}(\frac{2\pi e}{P}-k\pi)),&\mbox{if }e\leq E-\frac{kT}{2}\\ \tau^{+},&\mbox{elsewhere }\end{array}\right.$ (4)
where value $k\pi$ represents the delay of the cosine wave with respect to the starting epoch, $P$ is the period of the wave (Figure 2) and $k$ can be from $[0,1]$. $k=0$ reduces (4) to the regular cosine TS (CosSch), proposed by Kukleva et al. (2023) for the self-supervised tasks on the long-tailed datasets. With CosSch, the model starts training with a higher temperature $\tau^{+}$ and goes to a lower temperature $\tau^{-}$.
$\text{CosSch}(e;\tau^{+},\tau^{-},P)=\text{GCosSch}(e;\tau^{+},\tau^{-},P,k=0)$ (5)
However, we find that rather than using (5), it is beneficial for the task switching if we start with a lower temperature $\tau^{-}$ and move towards a higher value $\tau^{+}$. Starting with a lower temperature, the model provides priority to fewer neighbors, learning the coarse structure of representation space, resulting in a sharper decision boundary. The open set samples remain distant from any cluster of the known classes (due to heavy penalization of slight dissimilarity). As the temperature increases, the model prioritizes more neighbors and gradually pulls the positive samples to its own cluster. This makes within-class representations more compact and the decision boundary smoother while the core separation learned earlier is maintained. The open set samples are not pulled as tightly because the features are unknown to the model, maintaining the separation. This leads to a richer and potentially more generalized representation space for both open set and closed set performance. Figure 1 illustrates the concept with UMAP projection of representation spaces for different temperature schedules.
For $k=1$ in (4), the temperature starts with a lower value and goes to a higher one, looking like a negative cosine wave, hence the name negative cosine scheduling (NegCosSch).
$\text{NegCosSch}(e;\tau^{+},\tau^{-},P)=\text{GCosSch}(e;\tau^{+},\tau^{-},P,k=1)$ (6)
Through experimental results, we find that NegCosSch is better for OSR than CosSch or any other value of $k$ in GCosSch. The model settles down better if the final few epochs maintain a higher temperature rather than follow the wave (epochs 500-600 in Figure 2). Although initially aimed for OSR, the NegCosSch also improves performance for closed set classification while incurring no additional computational burden as it only includes an epoch-dependent temperature in a loss function.
3.3 Inference
For CE loss, we use the model as is for inference. However for SupCon loss, we remove the projection layer and a linear classifier is trained for evaluation. We use the maximum logit based scoring rule for OSR score.
4 Results and Discussion
In this Section, we describe the benchmarks used for evaluating our method, the experiment settings followed by results and discussion.
Benchmarks: We test our method on four regular OSR benchmarks: CIFAR10, CIFAR+10, CIFAR+50, and TinyImageNet. Moreover, we test our method on the SSBs defined on three fine-grained datasets: the Caltech-UCSD-Birds (CUB) (Wah et al. (2011)), FGVC-Aircraft (Krause et al. (2013)) and Stanford Cars (SCars) (Maji et al. (2013)). For these benchmarks, the open set classes are divided into ‘Easy’ and ‘Hard’ splits by computing the semantic similarity (based on the labeled visual attributes) of each pair of classes, the details of which can be found in Vaze et al. (2022). The different difficulty level along with more training classes make these datasets harder OSR benchmarks than the regular ones.
(a) CIFAR10
(b) CIFAR+10
(c) CIFAR+50
(d) TinyImageNet
Figure 3: Open Set AUROC of different temperature scheduling for the SupCon loss on the regular OSR benchmarks.
Training Details: We mostly follow the experiment settings and design choices from Vaze et al. (2022). For the regular benchmarks, we use a VGG32-like model and for the SSB datasets, we use a ResNet50 model pretrained on the places365 dataset [^2] In spite of our efforts, we could not find the same pretrained model mentioned in Vaze et al. (2022) online. Therefore, we use the pretrained model from Zhou et al. (2017) trained on places365, which is completely unrelated to the SSBs..
For experiments, we choose temperatures from $\mathcal{T}_{SupCon}=\{0.025,0.05,0.1,0.2,0.3,0.4,0.5\}$ for the SupCon loss and from $\mathcal{T}_{CE}=\{0.5,1.0,2.0\}$ for the CE loss. For label smoothing (LS) in SupCon loss, we implement the following function instead of (2):
$L_{\text{SC,LS}}=-\frac{1}{|I|}\sum_{i\in I}\sum_{j\in I\setminus\{i\}}\frac{1}{N_{i}(\alpha)}[(1-\alpha)\mathbf{1}_{y_{i}=y_{j}}+\frac{\alpha}{C-1}\mathbf{1}_{y_{i}\neq y_{j}}]\Bigg{[}\text{log}\frac{\text{exp}(\text{sim}(z_{i},z_{j})/\tau)}{\sum_{a\in I\setminus\{i\}}\text{exp}(\text{sim}(z_{i},z_{a})/\tau)}\Bigg{]}$ (7)
with $N_{i}(\alpha)=\sum_{k\in I\setminus\{i\}}[(1-\alpha)\mathbf{1}_{y_{i}=y_{k}}+\frac{\alpha}{C-1}\mathbf{1}_{y_{i}\neq y_{k}}]$, where $\alpha$ is the smoothing coefficient and $\mathbf{1}$ is the indicator function. The training details are provided in the Appendix. All the experiments are performed on a single NVIDIA RTX2080Ti GPU requiring from 2 to 32 hours to train a single model depending on the model and the dataset.
Metrics: We report the closed set performance as a $C$-class classification using accuracy (%), the open set performance as known-unknown detection using AUROC (%), and the open set classification rate (OSCR %). OSCR measures the trade-off between the open set and the closed set performance (Dhamija et al. (2018)). Now, we discuss the results.
4.1 NegCosSch yields better OSR performance (AUROC) than both CosSch and constant temperatures
Figure 3 presents the OSR performance of our proposed TS scheme along with the regular cosine scheduling and constant temperatures in SupCon loss on the regular benchmarks. We vary $(\tau^{+},\tau^{-})$ from $\mathcal{T}^{2}_{SupCon}$ and compare $\text{NegCosSch}(\tau^{+},\tau^{-},P)$ with $\text{CosSch}(\tau^{+},\tau^{-},P)$, $\text{Const}(\tau=\text{nearest}(\frac{1}{2}[\tau^{+}+\tau^{-}]))$ [^3] By ‘nearest’, we mean a nearest temperature is chosen from $\mathcal{T}_{SupCon}$. and $\text{Const}(\tau=\tau^{-})$. The objective of the quadruplet-wise comparisons is to determine if our proposed scheduling method outperforms a regular cosine scheduling, a constant temperature set to the midpoint of $(\tau^{+},\tau^{-})$, or set to $\tau^{-}$. We observe that for the CIFAR10, CIFAR+10 and TinyImageNet benchmarks, our proposed scheduling yields a better open set AUROC than CosSch and constant temperatures for most of the quadruplet comparisons. We find the improvement or degradation to be insignificant for the CIFAR+50 benchmark with the highest AUROC found for CosSch($0.3,0.1$).
Specifically, for any good value of constant temperature $\tau$ (which can be chosen from hyperparameter tuning) in the SupCon loss, it is better to use NegCosSch by setting $\tau^{+}=\tau+\Delta,\tau^{-}=\tau-\Delta$ (or alternatively, $\tau^{+}=\tau+\Delta,\tau^{-}=\tau$) with the increment Δ0.1 or $\Delta\approx 0.1\text{~{}or~{}}0.2$. For example, in TinyImageNet, we find from hyperparameter tuning that Const($\tau=0.2$) provides a good OSR performance. Therefore, NegCosSch($\tau^{+}=0.4,\tau^{-}=0.1$) or NegCosSch($\tau^{+}=0.3,\tau^{-}=0.2$) are better choices than Const($\tau=0.2$), CosSch($\tau^{+}=0.4,\tau^{-}=0.1$) and CosSch($\tau^{+}=0.3,\tau^{-}=0.2$). Utilizing this, we reduce the search space of $(\tau^{+},\tau^{-})$ on the SSB experiments, where we find higher performance boost discussed in Section 4.3. The closed set performance and AUROC with other values of $k$ in GCosSch are listed in the Appendix.
4.2 Improving OSR Methods with Our Proposed Scheme
Here to evaluate our scheme, we incorporate NegCosSch into three OSR methods: the CE baseline by Vaze et al. (2022), the SupCon baseline by Khosla et al. (2020) and the ConOSR by Xu et al. (2023) on the regular benchmarks. We improve the baseline scores of ConOSR for TinyImageNet with better design choices. The results shown in Table 1 indicate that adding our TS on top of these methods results in a similar or higher open set AUROC for most cases. We suspect that the benefits of our TS reduce when the number of training classes is small. For example, there are only 6 training classes in CIFAR10 and 4 training classes in the CIFAR+10 and CIFAR+50 benchmarks. For TinyImageNet with 20 training classes, our scheme shows significant improvements for all methods. We now consider the SSBs, where they have a larger number of training classes.
Table 1: Open set performance (AUROC %) comparison of different OSR methods with and without our proposed scheme.
Methods CIFAR10 CIFAR+10 CIFAR+50 TinyImageNet
CE (Baseline) (Vaze et al. (2022)) 93.6 97.9 96.5 83
CE + NegCosSch (ours) 93.76 97.88 96.32 83.47
SupCon (Baseline) 94.04 98.25 97.03 82.86
SupCon + NegCosSch (ours) 94.10 98.28 97.03 83.09
ConOSR (Xu et al. (2023)) 94.55 98.22 97.04 83.40
ConOSR + NegCosSch (ours) 94.52 98.29 97.01 83.63
Table 2: Performance on the SSBs. Open set results are shown on ‘Easy / Hard’ splits. Here, NC-TS means our negative cosine temperature scheduling and w/o LS means without label smoothing. We highlight the cases where our scheduling outperforms the corresponding baseline.
CUB FGVC-Aircraft SCars
Methods Acc. AUROC OSCR Acc. AUROC OSCR Acc. AUROC OSCR
ARPL (Chen et al. (2021)) 85.9 83.5 / 75.5 76.0 / 69.6 91.5 87.0 / 77.7 83.3 / 74.9 96.9 94.8 / 83.6 92.8 / 82.3
ARPL + NC-TS(ours) 85.97 87.46 / 79.5 79.38 / 73.5 90.58 91.06 / 81.40 85.76 / 77.46 96.97 95.82 / 86.24 93.83 / 84.95
CE (w/o LS) 84.60 83.40 / 74.68 70.59 / 63.29 90.29 91.33 / 81.75 82.38 / 73.95 96.66 93.20 / 84.32 90.17 / 81.70
CE (w/o LS) + NC-TS(ours) 86.20 86.96 / 77.67 74.91 / 67.01 90.95 91.82 / 83.33 83.57 / 76.1 97.49 94.59 / 86.22 92.25 / 84.14
CE + LS (Vaze et al. (2022)) 85.71 85.15 / 77.19 72.93 / 66.19 90.65 86.99 / 78.85 78.94 / 71.72 96.96 94.33 / 84.00 91.54 / 81.62
CE + LS + NC-TS(ours) 86.34 86.86 / 77.86 74.92 / 67.29 91.19 87.48 / 79.86 79.84 / 73.07 97.11 94.50 / 84.02 91.82 / 81.75
SupCon (w/o LS) 83.22 86.50 / 73.64 71.93 / 61.34 89.63 85.33 / 81.03 76.74 / 73.05 96.88 93.00 / 82.77 90.19 / 80.32
SupCon(w/oLS) + NC-TS(ours) 83.95 87.35 / 74.60 73.38 / 62.84 88.49 86.56 / 81.23 76.84 / 72.36 96.66 93.16 / 82.82 90.14 / 80.19
SupCon + LS 83.88 86.30 / 73.17 72.28 / 61.38 88.73 86.94 / 81.40 77.42 / 72.68 96.58 93.22 / 82.88 90.10 / 80.16
SupCon+LS+ NC-TS(ours) 83.81 86.49 / 74.40 72.38 / 62.34 89.57 85.36 / 81.29 76.57 / 73.10 96.73 93.49 / 83.05 90.49 / 80.44
The results are taken from Vaze et al. (2022).
4.3 Performance on the Semantic Shift Benchmarks
Here, we report the performance on the SSBs for both the previous OSR methods and our scheme in Table 2. We find only three previous methods in literature that report results on the SSBs. We implement the methods by Chen et al. (2021) and Vaze et al. (2022) by including our TS. As LS has shown performance improvements on SSBs, we experiment both with and without uniform LS for the CE and SupCon losses. Here, we do not optimize performance for $\alpha$ and temperatures but use a fixed set of hyperparameters for consistency. By following the strategy mentioned in Section 4.1, we fix $(\tau^{+},\tau^{-})$ at $(0.3,0.1)$ for TS and $\tau=0.2$ for reporting with constant temperatures in the SupCon loss. For CE loss, with the base temperature being $1.0$, we set $(\tau^{+},\tau^{-})$ at $(0.5,2.0)$.
We observe that if we include the proposed NegCosSch using either loss function, we get improved performances both for the closed set and open set problems over the corresponding baselines for all three datasets except for only the case of Aircraft on SupCon loss and including LS. Our NegCosSch provides performance boost for up to 1.60% of accuracy, up to 3.56%/2.99% of open set AUROC in the ‘Easy’/‘Hard’ splits and up to 4.32%/3.72% of OSCR. Moreover, our method can be used together with LS to further boost the performance in a few cases. This amount of performance boost comes without any additional computational cost.
For cases where LS does not help the open set performance (for the Stanford cars -‘Hard’ split and Aircraft with CE loss), our NegCosSch outperforms the baselines. Overall for SSBs, the CE loss performs better than SupCon loss. Comparison of these methods with CosSch is listed in the Appendix. Moreover, our method outperforms the OSR methods by Chen et al. (2021) and Vaze et al. (2022) in most cases. We believe that our TS scheme in principle can improve other OSR methods, such as the method by Jia et al. (2024).
(a) CUB
(b) SCars
(c) FGVC-Aircraft
Figure 4: Effect on performance improvement for our proposed scheduling over the baselines with varying number of training classes. Increasing the number of training classes tends to yield greater improvements in OSCR across all datasets, along with significant improvements in AUROC and accuracy, with the effect being most pronounced for CUB and least for FGVC-Aircraft. Error bars represent the standard deviation.
4.4 Proposed scheduling is more beneficial with higher number of training classes
To show the strength of our scheme with an increased number of training classes, we train models on the CE loss with $\{20\%,40\%,60\%,80\%\}$ of the randomly chosen training classes for the SSBs without changing the open set. In Figure 4, we plot the improvements (averaged across 5 trials) over the corresponding baseline. We define improvement of any metric $m$ as
$\text{improvement}=m[\text{NegCosSch}(\tau^{+},\tau^{-})]-\max_{\tau\in\mathcal{T}_{CE}}\{m[\text{Const}(\tau)]\}$
We observe an overall upward improvement trend for most cases in both ‘Easy’ and ‘Hard’ OSR splits in all three datasets as the number of training classes increases. The negative values in Figure 4(c) is due to the fact that we compare our method with the maximum scores of three baselines. With more training classes, the task becomes harder for the base model, which is observed by the performance decline. Nonetheless, our scheduling scheme gains higher improvement through TS.
4.5 Limitations
The open set performance depends on the nature of the unknown classes and their semantic similarity with the known classes. We find that the improvements over the baselines decrease, with a slight drop in closed set accuracy, when there is a small number of training classes, which occur in the CIFAR10, CIFAR+10, and CIFAR+50 benchmarks.
5 Related Works
Open Set Recognition: Since the introduction of OSR problem, it has received a significant interest in the research community. Most of the research attempts can be summarized into several common categories. The early ones are based on extreme value theory that modify the softmax function to model the open set samples (Vignotto and Engelke (2018); Bendale and Boult (2016)). Several works use generative models to generate synthetic unknown or counter-factual samples (Ge et al. (2017); Neal et al. (2018); Chen et al. (2021); Moon et al. (2022)), while other add auxiliary samples in the dataset with strategies, such as the sample mix-up (Chen et al. (2021); Xu et al. (2023); Bahavan et al. (2025); Li et al. (2024)), manifold mix-up (Zhou et al. (2021)), or by adding extra augmented samples (Wang et al. (2025); Jiang et al. (2023); Jia et al. (2024); Xu and Keuper ). Another group of research depends on training an additional model for a secondary objective function (Oza and Patel (2019); Sun et al. (2020); Perera et al. (2020); Yoshihashi et al. (2019); Zhang et al. (2020); Jia et al. (2024); Zhou et al. (2024)), for example training VAEs for computing the reconstruction loss hoping that it can improves the OSR score. However, training generative models or a secondary model is tough for larger benchmarks as it requires significant computation overhead.
Some methods either construct a different loss function (Chen et al. (2021, 2020a); Wang et al. (2022)) or add regularization to bound the open set risks (Zhou et al. (2021); Lu et al. (2022); Yang et al. (2024a)). For example, the method by Zhou et al. (2021) learns additional place-holders for the novel classes. Methods by Chen et al. (2021, 2020a) learn the reciprocal points of known classes signifying the representations of the ‘otherness’ corresponding to each class. These methods try to create additional empty regions in the representation space hoping that open set representations lie in those regions. The paper (Vaze et al. (2022)) shows that a well-trained closed set classifier is essential for good OSR. This has triggered research for better representation learning. For example, methods by Xu et al. (2023); Xu (2024); Li et al. (2024); Bahavan et al. (2025); Li et al. (2025) train models using the contrastive loss with regularization and heavy augmentations (Wang et al. (2025); Jiang et al. (2023); Jia et al. (2024)). The method by Wang et al. (2024) trains multiple experts for extracting diverse representations, and Yang et al. (2024b) proposes an open set self-learning framework, which adapts the model according to the test data assuming that it is available.
Temperature Scaling: Temperature scaling in the CE loss plays a crucial role in knowledge distillation (Hinton et al. (2015)), model calibration (Guo et al. (2017)) and so on. Since the contrastive loss has become popular for many tasks (Chen et al. (2020b); Khosla et al. (2020)), studies have aimed to understand its behavior. Works in Wang and Liu (2021); Kukleva et al. (2023); Zhang et al. (2021, 2022) explain the role of temperature in self-supervised contrastive learning. Recently, methods in Jeong et al. (2024); Qiu et al. (2024) utilize temperature cool-down methods in language models.
6 Conclusion
We develop a novel temperature scheduling scheme, which can be folded into any existing OSR method with both cross-entropy and contrastive losses without any computational overhead. We find that starting with a lower temperature and moving towards a higher temperature results in making tighter representation clusters for the closed set classes, while the representations of the open set examples remain more distant. This process is more effective than using a fixed temperature or the opposite temperature scheduling scheme. We demonstrate improvement of open set performance for some of the well-known OSR methods on the regular benchmarks when we include our temperature scheduling. Our proposed scheduling demonstrates stronger performance improvements for the tougher semantic shift benchmarks both for the closed set and the open set problems. The benefit of our scheme scheme can be better realized with a larger number of training classes.
7 Acknowledgments
We are thankful to the University of Maryland Institute for Health Computing for their IT support throughout the project.